You order some accessories online and get an estimated delivery date of June 1–June 11. You know you will be out of town June 4th, 5th, and 6th and are a little concerned about the package arriving when you are away. Assuming the delivery date follows a discrete uniform distribution, what is the likelihood your package will be delivered while you are out of town? Round your answer to four decimal places, if necessary.
1
Expert's answer
2020-10-01T15:11:56-0400

Solution

X~ discrete uniform (1,11)


"f(x) = {1 \over b-a+1}""f(x) = {1 \over 11-1+1}={1 \over 11}"

The probability of the package arriving on any of the days is 1/11.

The number of days you'll be out of town is 3

Thus we calculate P(X=3)


"F(x) =P(X \le x) ={ x-a+1 \over b-a+1}"

"={ x-1+1 \over 11-1+1}={x\over 11}"


"P(X \le 3)={3 \over 11} = 0.2727"

Answer : 3/11 or 0.2727